OpenStudy (anonymous):

solve using the fundamental of calculus integrate [1,4] sqrt(5/x)dx

OpenStudy (anonymous):

idk why i have a diffirent numbers one is \[3\sqrt{5} \]

OpenStudy (anonymous):

and the other is \[2 \sqrt{5}\]

OpenStudy (anonymous):

i mean different answer

OpenStudy (anonymous):

\[\sqrt{\frac{5}{x}} \]What is the integral of this? can you tell me what you get?

OpenStudy (anonymous):



OpenStudy (anonymous):